OpenStudy (laila_2278):

,What is the factored form of the expression? 10x^2 – 43xy – 66y^2 5. x^2 – 10xy + 24y^2

OpenStudy (anonymous):

I would factor by grouping. \[10x^2 – 43xy – 66y^2 \] The coefficient of x^2 and y^2 is 10 and -66. The product of that is \[10*(-66)=-660\] The factors of -660 which sum to -43 (comes form the coefficient of xy.) is -55 and 12. \[-55*12=-660\] So you get: \[10x^2 – 43xy – 66y^2=10x^2-55xy+12xy-66y^2 \] \[=2x(5x+6y)-11y(5x+6y)\] Here you can factor common terms: \[(5x+6y)(2x-11y)\]

OpenStudy (laila_2278):

14. Find the area of the shaded region. Use the formula A = The term is pi.r2 to find the area of a circle. https://www.connexus.com/content/media/666994-8212014-91604-PM-1020085866.jpg